# How Does a Virtual Agent Decide Where to Look? - Symbolic Cognitive Reasoning for Embodied Head Rotation

### v1

@@ Abstract
Figure 1: This example illustrates how SCORE produces human-like head movements in a multi-agent setting. The highlighted agent attends to the intended walking path, aligning closely with the ground-truth behavior. In contrast, Track fixates excessively on static background elements (e.g., storefronts), failing to handle multiple factors (e.g., secure path, understand pedestrian flow) and resulting in implausible head rotations.
1 Introduction
Realistic and context-aware virtual agents have long been a central research focus in computer graphics [1, 2, 3, 4]. One of the key aspects of believable agent behavior lies in natural head rotations that reveal what it notices, what it intends, and how it handles risk.
Synthesizing such micro-behavior is a long-standing challenge. Yet recent head movement synthesizers still treat attention as a colored blob on the image plane. Trained on saliency maps [5, 6, 7, 8], they steer the camera toward the conspicuous elements while ignoring why a human might look elsewhere. Real pedestrians, for example, repeatedly scan the flow of oncoming cars, bicycles, and walkers, balancing safety, curiosity, and social convention in a single sweep. Saliency-only models over-predict attraction to bright or moving objects and under-predict purposeful reorientations guided by cognitive processes, degrading realism in cluttered, dynamic scenes.
@@ 1 Introduction
2 RELATED WORK
2.1 Head-Movement Prediction
Head orientation is widely regarded as a reliable proxy for observable visual attention: in social activity or outdoor tasks, where a person turns their head often signals what they are thinking about and concentrating on [9, 10, 11, 12]. In other words, predicting those turns requires more than stimulus detection; it calls for an understanding of the underlying cognitive intent.
Head movement prediction has been widely researched to optimize user experience in 360-degree video consumption [13, 14]. Early methods extrapolated head trajectories via linear filters [15, 16], but they ignored the link between attention and changing content. Subsequent work fused saliency with motion history; PanoSalNet + LSTM captured the equatorial bias of HMD viewers and improved accuracy during rapid turns [12]. TRACK advanced this line by balancing trajectory and saliency cues with a Structural-RNN [17], achieving state-of-the-art robustness across exploratory and focus-driven clips [8].
Despite these gains, saliency- and trajectory-based predictors remain stimulus-bound; they overlook the cognitive drivers—information seeking, habitual scanning—that govern human re-orientation. [4] argue that physically plausible motion is tied to higher-level reasoning. Our work follows this direction, using symbolic cognitive predicates and vision–language reasoning to predict not just where an agent looks, but why.
2.2 Vison-Language Reasoning
Recent advancements in VLM and LLM have enabled their application across various domains, including scene understanding [18], mathematical problem solving [19], and embodied agent control [20]. They increasingly rely on Chain-of-Thought prompting [21], which forces models to spell out intermediate reasoning  [22, 23]. By inserting an explicit reasoning layer between perception and action, these systems plan more reliably and generalise across tasks [24, 25].
@@ 2.2 Vison-Language Reasoning
3 Motivation and Problem Statement
Research on realistic virtual-agent behavior has traditionally focused on macro-level phenomena, such as crowd simulations [26, 27, 28] or trajectory prediction [29, 30, 31, 32, 33]. By comparison, micro-level actions such as head-orientation remain under-explored, even though subtle head shifts strongly influence perceived plausibility in interactive VR and game settings.
Synthesizing natural head turns is challenging because the motion arises from the integration of scene perception, cognitive evaluation, and decision-making. Rather than explicitly modeling these intricate processes, recent studies typically adopt data-driven approaches or visual-saliency proxies [8, 12]. Although such approaches capture conspicuous cues, their shallow treatment of cognition limits generalizability: agents fail to adapt when salient features conflict with task context or social norms, leading to perceptually implausible behaviour in environments.
@@ 4 User Study
Figure 2: Red dashed boxes highlight elements present only in the APC version of each environment: a Santa, a traffic accident, and arguing people. The corresponding MDC scenes omit such distracting events.
Figure 3: Categorized distribution of participants’ self-reported head-movement rationales.
4.1 Apparatus and Settings
@@
4.2 Method and Procedure
A user study was conducted with two experimenters. Upon arrival, each participant provided informed consent and demographic data, then completed a 15-minute orientation (5-minute briefing and leq\ leq10-minute free navigation in a neutral city scene). Simulator Sickness Questionnaire [34] (SSQ) scores taken before and after the study showed no significant change, and Simulation Task Load Index [35] (SIM-TLX) ratings confirmed low mental workload.
@@ 4.2 Method and Procedure
We recorded egocentric video, head-orientation quaternions, and sensor data during each scenario. Upon completion of all scenarios, participants reviewed their footage and articulated the underlying motive for every head turn. To facilitate reliable articulation, participants were presented with example prompts such as “it caught my eye," or “I needed to confirm something.
4.3 Derivation of Motivational Drivers
@@
We annotated each trajectory and its 200 accompanying participant explanations. This began with a codebook containing the three motivational factors derived from the attention research—Safety, Interest, and Information Seeking [36, 37, 38]. During iterative review, we observed recurrent behaviours that these labels could not capture, such as surveying the general flow of pedestrians in a shopping mall or tilting the head upward to scan the sky when walking. This led us to define two additional categories, Social Schema and Habit. The resulting five-factor set constitutes the motivational taxonomy that supports SCORE’s decision model. Figure 3 visualises how often each motivation category appeared in participant explanations across the various scenarios.
Figure 4: SCORE pipeline. The two-stage architecture combines offline deliberation with near-online refinement. In the DPS, PEM passes each “novel” view, the entity lists, and the scenario goal to a VLM, which generates contextual descriptions stored in FMM. The Planning Module then invokes an LLM that reasons over the five cognitive drivers to select an action–reason pair that schedules the next head orientation. During runtime, RES evaluates a 2-second look-ahead image with a lightweight FastVLM; if the pre-planned action is inconsistent with the live context, it is replaced before execution.
5 SCORE Framework
SCORE operates through a two-stage pipeline. Deliberative Perception–Planning Stage (DPS) is the offline planning stage that predicts and schedules the agent’s desired head orientation along the agent’s trajectory. Reactive Execution Stage (RES) is the near-online validation stage that refines the pre-planned desired head orientation during runtime. The overall pipeline is shown in Figure 4; implementation details appear in the supplementary.
5.1 Deliberative Perception–Planning Stage (DPS)
DPS runs once for an entire trajectory. At every 0.2s, it submits the agent’s current field of view (FOV) to the Perception Module (PEM). To avoid redundant processing, PEM compares the incoming view image with the most recent “novel" view and flags it as novel only if the Structural Similarity Index Measure (SSIM) [39] between them is below 60 [40]. For each novel view, a forward ray-tracing pass enumerates visible objects and agents. The novel view image, the resulting entity lists, and the scenario goal are then passed to the VLM, which produces contextual descriptions for every entity. These lists and descriptions, stamped with the capture time, are stored in the Foundational Memory Module (FMM).
@@ 5.1 Deliberative Perception–Planning Stage (DPS)
DPS simulates the head turn toward the target at $36^{\circ}/sec$ [41]. During the turn, PEM is paused. Once the target pose is achieved, the agent holds its head for 1-2 seconds $\sim N(1.5,0.25^{2})$. The head then returns to the walking direction, mimicking the human tendency to realign the head after a temporal goal (e.g., threat assessment or curiosity) has been met. PEM resumes after the first half of the head’s hold period, ready to identify the next novel view.
Each FMM entry contains the time stamp, the current object and agent lists, and the associated action–reason pair. To maintain robust performance over long horizons, the FMM retains at most twenty entries: ten are the most recent, and the remainder are those deemed most relevant to the scenario goal according to an LLM-based relevance SCORE, following recent agent-memory strategies [42].
@@
5.2 Reactive Execution Stage (RES)
RES operates continuously during the simulation. Every 0.2s, it assembles the agent’s current FOV, self-pose, entity list, the scenario goal, the most recent action-reason pair, and the list of actions that actually been performed, and submits this bundle to a lightweight FastVLM. FastVLM determines whether the pre-planned pose remains valid or should be revised. Executed action is logged via the manage schema used in FMM for ensuring dynamic adaptability.
@@ 5.2 Reactive Execution Stage (RES)
6 Evaluation
In our evaluation, we use Gemini 1.5 Pro as the VLM and LLM for DPS, and Gemini 1.5 Flash [43] as the FastVLM for RES. To assess SCORE’s ability to replicate human-like head-motion patterns, we benchmark it against the saliency-driven baseline Track [8] and the user-generated motion data collected in the user study (hereafter, Human). Similar to prior work that classifies nods and shakes from orientation traces [44, 45], we employ Dynamic Time Warping (DTW) [46] to compare each simulated quaternion sequence with its human counterpart. DTW is well suited to this task: it aligns two time series sample-by-sample while accommodating local speed variations, yielding an interpretable measure of how closely a generated trajectory follows the spatial path taken by real users [47, 48].
Table 1: DTW comparison between IntraHuman, Track, and SCORE in single-agent settings across five scenarios under MDC and APC conditions. Lower DTW scores indicate closer alignment with human motion data. The best-performing scores are highlighted in bold. The reported $\pm$ values denote 95% confidence intervals.
@@ 6 Evaluation
Table 2: DTW comparison between Track, and SCORE in multi-agent settings. The best-performing scores are highlighted in bold. The reported $\pm$ values denote 95% confidence intervals.
@@
6.1 Single-Agent Benchmark
6.1.1 Method
We evaluated SCORE on the five distinct virtual scenarios—Bus, Café, Crossing, Mall, and Street—under both MDC and APC conditions, as used in the user study. For each scenario–condition pair, we generated five head-motion trials with SCORE and with Track, using the same agent body trajectories that were recorded from the user study and replaying them at a constant walking speed of $1.3m/s$ [49]. All hyperparameters and VLM/LLM prompts were held fixed across scenarios, ensuring a fair comparison.
@@ 6.1.1 Method
$d(q_{1},q_{2})=2\cdot\arccos\left(\left|\text{dot}(q_{1},q_{2})\right|\right),$ (1)
@@ 6.1.2 Results
The advantage is most pronounced in the Mall and Crosswalk, where rapid scene changes and the presence of multiple objects. SCORE also excels in the Bus-APC condition: it consistently attends to a Santa Claus passenger who captured participants’ attention, showing that its cognitive layer can infer the social salience of atypical objects.
@@
6.2 Multi-Agent Benchmark
6.2.1 Method
To evaluate SCORE’s generalizability in multi-agent scenarios, we conducted an additional experiment on the UCY crowd dataset [50], which provides frame-level positions, head orientations, and corresponding video footage. We reproduced three scenes (zara01, zara02, and students003) in our simulator by replicating the recorded agent configurations and environmental layout. The model then generated head-motion trajectories under these conditions, and similarity to the ground-truth orientations was quantified with the DTW metric.
6.2.2 Results
@@
Table 3: Responsiveness benchmark with unseen distractors. Any improvement of MDC-APC over MDC-APC reflects RES’s ability to detect and adapt to novel stimuli not available at planning time.
@@
Table 4: Ablation study of SCORE in single-agent settings across five scenarios under MDC and APC conditions.
@@
Table 5: Ablation study of SCORE in multi-agent settings across three scenes.
@@
Figure 5: Visualization of action and reasoning result when Interest is removed from Motivational Drivers.
6.3 Responsiveness to Unseen Distractors

### v2

# How Does a Virtual Agent Decide Where to Look? Symbolic Cognitive Reasoning for Embodied Head Rotation

@@ Abstract.
Virtual agents, head movements prediction, visual language reasoning
$c$s: Human-centered computing Virtual realityccs: Computing methodologies Intelligent agents
1. Introduction
Realistic and context-aware virtual agents have long been a central research focus in computer graphics (Lerner et al., 2010; Steel et al., 2010; Narang et al., 2016; Curtis et al., 2022). One of the key aspects of believable agent behavior lies in natural head rotations that reveal what it notices, what it intends, and how it handles risk.
Synthesizing such micro-behavior is a long-standing challenge. Yet recent head movement synthesizers still treat attention as a colored blob on the image plane. Trained on saliency maps (Xu et al., 2018; Zhu et al., 2020; Yang et al., 2021; Rondon et al., 2022), they steer the camera toward the conspicuous elements while ignoring why a human might look elsewhere. Real pedestrians, for example, repeatedly scan the flow of oncoming cars, bicycles, and walkers, balancing safety, curiosity, and social convention in a single sweep. Saliency-only models over-predict attraction to bright or moving objects and under-predict purposeful reorientations guided by cognitive processes, degrading realism in cluttered, dynamic scenes.
@@ 1. Introduction
2. RELATED WORK
2.1. Head-Movement Prediction
Head orientation is widely regarded as a reliable proxy for observable visual attention: in social activity or outdoor tasks, where a person turns their head often signals what they are thinking about and concentrating on (Stiefelhagen et al., 2002; Danninger et al., 2005; Ba and Odobez, 2008; Nguyen et al., 2018). In other words, predicting those turns requires more than stimulus detection; it calls for an understanding of the underlying cognitive intent.
Head movement prediction has been widely researched to optimize user experience in 360-degree video consumption (Fan et al., 2017; Aladagli et al., 2017). Early methods extrapolated head trajectories via linear filters (Qian et al., 2016; Duanmu et al., 2017), but they ignored the link between attention and changing content. Subsequent work fused saliency with motion history; PanoSalNet + LSTM captured the equatorial bias of HMD viewers and improved accuracy during rapid turns (Nguyen et al., 2018). TRACK advanced this line by balancing trajectory and saliency cues with a Structural-RNN (Jain et al., 2016), achieving state-of-the-art robustness across exploratory and focus-driven clips (Rondon et al., 2022).
Despite these gains, saliency- and trajectory-based predictors remain stimulus-bound; they overlook the cognitive drivers—information seeking, habitual scanning—that govern human re-orientation. This gap highlights the importance of believability: beyond mechanical accuracy, motion must express underlying intentions—what classic animation has described as the “illusion of life” (Thomas, 1995). In a similar spirit, (Curtis et al., 2022) argue that plausible motion is tied to higher-level reasoning. Our work follows this direction, using symbolic cognitive predicates and vision–language reasoning to predict not just where an agent looks, but why.
2.2. Vison-Language Reasoning
Recent advancements in VLM and LLM have enabled their application across various domains, including scene understanding (Fu et al., 2024), mathematical problem solving (Liang et al., 2023), and embodied agent control (Kim et al., 2024). They increasingly rely on Chain-of-Thought prompting (Wei et al., 2022), which forces models to spell out intermediate reasoning  (Yao et al., 2023; Shinn et al., 2023). By inserting an explicit reasoning layer between perception and action, these systems plan more reliably and generalise across tasks (Song et al., 2023; Yang et al., 2023).
@@ 2.2. Vison-Language Reasoning
3. Motivation and Problem Statement
Research on realistic virtual-agent behavior has traditionally focused on macro-level phenomena, such as crowd simulations (Panayiotou et al., 2022; Charalambous et al., 2023; Ji et al., 2024) or trajectory prediction (Yue et al., 2022; Guo et al., 2022; Lin et al., 2025; Wong et al., 2022; Mangalam et al., 2021). By comparison, micro-level actions such as head-orientation remain under-explored, even though subtle head shifts strongly influence perceived plausibility in interactive VR and game settings.
Synthesizing natural head turns is challenging because the motion arises from the integration of scene perception, cognitive evaluation, and decision-making. Rather than explicitly modeling these intricate processes, recent studies typically adopt data-driven approaches or visual-saliency proxies (Rondon et al., 2022; Nguyen et al., 2018). Although such approaches capture conspicuous cues, their shallow treatment of cognition limits generalizability: agents fail to adapt when salient features conflict with task context or social norms, leading to perceptually implausible behaviour in environments.
@@ 4. User Study
Figure 2. Red dashed boxes highlight elements present only in the APC version of each environment: a Santa, a traffic accident, and arguing people. The corresponding MDC scenes omit such distracting events.
4.1. Apparatus and Settings
@@
Figure 3. Categorized distribution of participants’ self-reported head-movement rationales.
Figure 4. SCORE pipeline. The two-stage architecture combines offline deliberation with near-online refinement. In the DPS, PEM passes each “novel” view, the entity lists, and the scenario goal to a VLM, which generates contextual descriptions stored in FMM. The Planning Module then invokes an LLM that reasons over the five cognitive drivers to select an action–reason pair that schedules the next head orientation. During runtime, RES evaluates a 2-second look-ahead image with a lightweight FastVLM; if the pre-planned action is inconsistent with the live context, it is replaced before execution.
4.2. Method and Procedure
A user study was conducted with two experimenters. Upon arrival, each participant provided informed consent and demographic data, then completed a 15-minute orientation (5-minute briefing and up to 10 minutes of free navigation in a neutral city scene). Simulator Sickness Questionnaire (Kennedy et al., 1993) (SSQ) scores taken before and after the study showed no significant change, and Simulation Task Load Index (Harris et al., 2020) (SIM-TLX) ratings confirmed low mental workload.
@@ 4.2. Method and Procedure
We recorded egocentric video, head-orientation quaternions, and sensor data during each scenario. Upon completion of all scenarios, participants reviewed their footage and articulated the underlying motive for every head turn. To facilitate reliable articulation, participants were presented with example prompts such as “it caught my eye,” or “I needed to confirm something.
4.3. Derivation of Motivational Drivers
@@
We annotated each trajectory and its 200 accompanying participant explanations. This began with a codebook containing the three motivational factors derived from the attention research—Safety, Interest, and Information Seeking (Yang et al., 2024; Witchel et al., 2016; Bromberg-Martin and Monosov, 2020). During iterative review, we observed recurrent behaviours that these labels could not capture, such as surveying the general flow of pedestrians in a shopping mall or tilting the head upward to scan the sky when walking. This led us to define two additional categories, Social Schema and Habit. The resulting five-factor set constitutes the motivational taxonomy that supports SCORE’s decision model. Figure 3 visualises how often each motivation category appeared in participant explanations across the various scenarios.
5. SCORE Framework
SCORE operates through a two-stage pipeline. Deliberative Perception–Planning Stage is the offline planning stage that predicts and schedules the agent’s desired head orientation along the agent’s trajectory. Reactive Execution Stage is the near-online validation stage that refines the pre-planned desired head orientation during runtime. The overall pipeline is shown in Figure 4; implementation details appear in the supplementary.
5.1. Deliberative Perception–Planning Stage (DPS)
DPS runs once for an entire trajectory. At every 0.2s, it submits the agent’s current field of view (FOV) to the Perception Module (PEM). To avoid redundant processing, PEM compares the incoming view image with the most recent “novel” view and flags it as novel only if the Structural Similarity Index Measure (SSIM) (Wang et al., 2004) between them is below 60 (Yedla et al., 2020). For each novel view, a forward ray-tracing pass enumerates visible objects and agents. The novel view image, the resulting entity lists, and the scenario goal are then passed to the VLM, which produces contextual descriptions for every entity. These lists and descriptions, stamped with the capture time, are stored in the Foundational Memory Module (FMM).
@@ 5.1. Deliberative Perception–Planning Stage (DPS)
DPS simulates the head turn toward the target at $36^{\circ}/sec$ (Grossman et al., 1988). During the turn, PEM is paused. Once the target pose is achieved, the agent holds its head for 1-2 seconds $\sim N(1.5,0.25^{2})$. The head then returns to the walking direction, mimicking the human tendency to realign the head after a temporal goal (e.g., threat assessment or curiosity) has been met. PEM resumes after the first half of the head’s hold period, ready to identify the next novel view.
Each FMM entry contains the time stamp, the current object and agent lists, and the associated action–reason pair. To maintain robust performance over long horizons, the FMM retains at most twenty entries: ten are the most recent, and the remainder are those deemed most relevant to the scenario goal according to an LLM-based relevance SCORE, following recent agent-memory strategies (Seo et al., 2025).
@@
5.2. Reactive Execution Stage (RES)
RES operates continuously during the simulation. Every 0.2s, it assembles the agent’s current FOV, self-pose, entity list, the scenario goal, the most recent action-reason pair, and the list of actions that have actually been performed, and submits this bundle to a lightweight FastVLM. FastVLM determines whether the pre-planned pose remains valid or should be revised. Executed action is logged via the manage schema used in FMM for ensuring dynamic adaptability.
@@ 5.2. Reactive Execution Stage (RES)
6. Evaluation
In our evaluation, we use Gemini 1.5 Pro as the VLM and LLM for DPS, and Gemini 1.5 Flash (Team et al., 2024) as the FastVLM for RES. To assess SCORE’s ability to replicate human-like head-motion patterns, we benchmark it against the saliency-driven baseline Track (Rondon et al., 2022) and the user-generated motion data collected in the user study (hereafter, Human). Similar to prior work that classifies nods and shakes from orientation traces (Switonski et al., 2019; Li and Hu, 2024), we employ Dynamic Time Warping (DTW) (Sakoe and Chiba, 1978) to compare each simulated quaternion sequence with its human counterpart. DTW is well-suited to this task: it aligns two time series sample-by-sample while accommodating local speed variations, yielding an interpretable measure of how closely a generated trajectory follows the spatial path taken by real users (Cuturi and Blondel, 2017; Lerogeron et al., 2023). However, because human trajectories vary considerably across individuals even under identical scenarios, due to personal priorities and habits, no single trajectory can serve as a definitive ground truth. To address this, we compare each generated trajectory against all human trajectories collected for the same scenario and report the averaged DTW distance as a similarity score.
At the same time, DTW evaluates trajectory-level alignment but does not capture the perceived believability of the behavior. To complement this limitation, we conducted a small user study to assess realism and plausibility. The outcomes of this study are provided in the supplementary material.
Table 1. DTW comparison between IntraHuman, Track, and SCORE in single-agent settings across five scenarios under MDC and APC conditions. Lower DTW scores indicate closer alignment with human motion data. The best-performing scores are highlighted in bold. The reported $\pm$ values denote 95% confidence intervals.
@@ 6. Evaluation
Table 2. DTW comparison between Track, and SCORE in multi-agent settings. The best-performing scores are highlighted in bold. The reported $\pm$ values denote 95% confidence intervals.
@@
6.1. Single-Agent Benchmark
6.1.1. Method
We evaluated SCORE on the five distinct virtual scenarios—Bus, Café, Crossing, Mall, and Street—under both MDC and APC conditions, as used in the user study. For each scenario–condition pair, we generated five head-motion trials with SCORE and with Track, using the same agent body trajectories that were recorded from the user study and replaying them at a constant walking speed of $1.3m/s$ (Bohannon, 1997). All hyperparameters and VLM/LLM prompts were held fixed across scenarios, ensuring a fair comparison.
@@ 6.1.1. Method
(1) $d(q_{1},q_{2})=2\cdot\arccos\left(\left|\text{dot}(q_{1},q_{2})\right|\right),$
@@ 6.1.2. Results
The advantage is most pronounced in the Mall and Crosswalk, where rapid scene changes and the presence of multiple objects occur. SCORE also excels in the Bus-APC: it consistently attends to a Santa Claus passenger who captured participants’ attention, showing that its cognitive layer can infer the social salience of atypical objects.
@@
6.1.3. Categorical Analysis of Entity Selection
While DTW evaluation quantifies the similarity of rotational trajectories between humans and the model, it does not reveal whether both selected similar entities within the scene. To complement this, we compared the distribution of entity categories chosen by humans and our SCORE model in each scenario (Figure 5). Under the MDC, humans most frequently selected the forward direction, followed by pedestrians and traffic lights. SCORE also prioritized the forward direction but assigned relatively greater weight to traffic lights. We speculate that SCORE may exhibit a more “safety-conscious” behavior by periodically checking traffic lights to ensure safe navigation.
In contrast, under the APC, although the task remained the same, the newly introduced accident scene strongly attracted human attention, leading both humans and the model to allocate substantial attention to this entity. Humans tended to secure safety by scanning both left and right while crossing the street, whereas the agent focused more on the left—the side of the accident—when ensuring safety. Both strategies appear reasonable given the circumstances.
Collectively, the results suggest that both humans and the model turned their gaze for their own valid reasons. The observed differences are less indicative of outright mistakes than of distinct tendencies—or even persona-like variations—in how attention was allocated. See Figure 5 for the full distribution across scenes; detailed results for the bus,café, mall, and street are provided in the supplementary material.
Figure 5. The figure illustrates differences in proportional selections, which indicate prioritization tendencies between humans and the model in the MDC and APC of the crossing scene.
6.2. Multi-Agent Benchmark
6.2.1. Method
To evaluate SCORE’s generalizability in multi-agent scenarios, we conducted an additional experiment on the UCY crowd dataset (Lerner et al., 2007), which provides frame-level positions, head orientations, and corresponding video footage. We reproduced three scenes (zara01, zara02, and students003) in our simulator by replicating the recorded agent configurations and environmental layout. The model then generated head-motion trajectories under these conditions, and similarity to the ground-truth orientations was quantified with the DTW metric.
6.2.2. Results
@@
Table 3. Responsiveness benchmark with unseen distractors. Any improvement of MDC-APC over MDC-APC reflects RES’s ability to detect and adapt to novel stimuli not available at planning time.
@@
Table 4. Ablation study of SCORE in single-agent settings across five scenarios under MDC and APC conditions.
@@
Table 5. Ablation study of SCORE in multi-agent settings across three scenes.
@@
Figure 6. Visualization of action and reasoning result when Interest is removed from Motivational Drivers.
6.3. Responsiveness to Unseen Distractors